# The formation of a soliton gas condensate for the focusing Nonlinear Schrödinger equation

### v1

@@ 2.5 Riemann-Hilbert problem for the soliton gas
where we recall that the contour $\eta_{1}$ is oriented from right to left, while $\eta_{2}$ is oriented from left to right, and in both cases the error term $\mathcal{O}\left(\frac{1}{N}\right)$ is uniform for all $z$ in the respective contour $\Gamma_{1}$ or $\Gamma_{2}$. Indeed, under the $N$-soliton condensate assumption, a complete asymptotic expansion for this sum can be calculated, because the eigenvalues $\{\lambda_{j}\}_{j=1}^{N}$ have been selected to arrive at a uniform mid-point approximation. The next term in the expansion can be computed explicitly, we perform such computation in Appendix A.
@@
  • $\tilde{A}(z)=I+\frac{1}{2iz}\left(\begin{array}[]{cc}-\int_{x}^{\infty}|\psi_{SG}(s,t;N)|^{2}{\rm d}s&\psi_{SG}(x,t;N)\\ \overline{\psi_{SG}(x,t;N)}&\int_{x}^{\infty}|\psi_{SG}(s,t);N|^{2}{\rm d}s\end{array}\right)+\mathcal{O}\left(\frac{1}{z^{2}}\right),\ \mbox{ as }z\to\infty$.
@@
This permits the direct application of an existence theory via the connection between Riemann-Hilbert problems and Fredholm singular integral operators (see Theorems 9.1, 9.2, and 9.3 in [42], or the proof of Theorem 5.3 in [10]). Thus we state without further discussion the following result:
@@ 2.6 Statement of the results
Under the $N$-soliton condensate scattering data assumption, for all $(x,t)$ in a compact set $K$, there is $N_{0}$ so that for all $N>N_{0}$, $\psi_{SG}(x,t;N)$ satisfies the asymptotic description (7.8). Specifically, for $\left|\psi_{SG}\right|$, we have
$|\psi_{SG}(x,t;N)|^{2}=c^{2}\frac{\left(\Theta^{\prime}\left(\frac{\tau+1}{2};\tau\right)\right)^{2}}{\Theta(\tau;\tau)\Theta(0;\tau)}\frac{\Theta\left(-\frac{\zeta(x,N)}{2\pi}+\frac{1-\tau}{2};\tau\right)\Theta\left(-\frac{\zeta(x,N)}{2\pi}+\frac{\tau-1}{2};\tau\right)}{\Theta^{2}\left(-\frac{\zeta(x,N)}{2\pi};\tau\right)}\ +\ \mathcal{O}\left(\frac{1}{\log{N}}\right),$ (2.36)
where the error term $\mathcal{O}\left(\frac{1}{\log{N}}\right)$ is uniform for all $(x,t)$ in the compact set $K$. Here $\Theta(z;\tau)$ is the Jacobi Theta-3 function as defined in [12]. The constants $\tau,c$ and $\zeta(x,N)$ are given by
$c=\frac{1}{2}\left(\int_{A}^{-{\overline{A}}}\frac{1}{R_{+}(s)}{\rm d}s\right)^{-1}\,,\quad\tau=2c\int_{-{\overline{A}}}^{-A}\frac{1}{R(s)}{\rm d}s\,,$ (2.37)
$\zeta(x,N)=-\frac{\tau\Delta(x)-2\ln(N)}{2c\pi}\left(\int_{-{\overline{A}}}^{-A}\frac{s^{2}}{R(s)}{\rm d}s-\tau\int_{A}^{-{\overline{A}}}\frac{s^{2}}{R_{+}(s)}{\rm d}s\right),$ (2.38)
where $R(z)=(z-A)^{\frac{1}{2}}(z+A)^{\frac{1}{2}}(z-{\overline{A}})^{\frac{1}{2}}(z+{\overline{A}})^{\frac{1}{2}}$, the brunch-cuts are the canonical ones, and the integration is performed on the straight line connecting the endpoints. The function $\Delta(x)$ is defined in terms of the functions $\rho(z)$, $h(z)$ as follows
$\Delta(x)=\left(\int_{-A}^{-{\overline{A}}}\frac{1}{R(s)}{\rm d}s\right)^{-1}\left(-\int_{A}^{-{\overline{A}}}\frac{\log\,(2\pi h(s)\rho(s))+2isx}{R_{+}(s)}{\rm d}s+\int_{-A}^{{\overline{A}}}\frac{\log\left(2\pi h(s)\rho(s)\right)-2isx}{R_{+}(s)}{\rm d}s\right)\,.$ (2.39)
@@
In appendix B we consider the non-symmetric setting where the line segments $\eta_{1}$ and $\eta_{2}$ are shifted so that they are not symmetric with respect to the origin. An argument involving a translation of the spectral plane reveals a new explicit dependence on time. Finally, it is worth noticing that the solution depends just on the product of $h(z),\rho(z)$ and not on each function separately. Fig. 3 shows plots of the leading order asymptotic of the solution of the NLS equation for different values of $N$.
@@
Let $A(z),{\widetilde{A}}(z)$ be the solutions of RHP 2.3 and RHP 2.4 respectively. Under the $N$-soliton condensate scattering data assumption, there is $N_{0}$ so that for all $N>N_{0}$, we have
$A(z)=\left(I_{2}+\mathcal{O}\left(\frac{1}{N\left(1+|z|\right)}\right)\right){\widetilde{A}}(z)\,.$ (2.40)
where the error term, $\mathcal{O}\left(\frac{1}{N\left(1+|z|\right)}\right)$ is uniform in the entire complex plane.
@@
In the next four sections, we develop the asymptotic analysis of the solution of the Riemann-Hilbert problem 2.4, to establish, in Section 7, the asymptotic behavior of the soliton gas condensate solution $\psi_{SG}(x,t;N)$. We prove Proposition 2.7, Corollary 2.8 and other technical results in Appendix A.
@@ 3 Dressing transformations
$C(z)=e^{(\ln(N)g(\infty)+y(\infty))\sigma_{3}}B(z)e^{-(\ln(N)g(z)+y(z))\sigma_{3}}\,,$ (3.3)
where $g(\infty),y(\infty)$ are the values of $g(z)$ and $y(z)$ at $\infty$, which we assume are finite (this assumption will be verified later, since these two quantities will be explicitly constructed). The function $C(z)$ satisfies the following RHP.
@@
    $g_{+}(z)-g_{-}(z)={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{2}{\tau}}},\quad z\in\gamma$ (3.6)
  • $g(z)=g(\infty)+\mathcal{O}(1/z)$, as $z\to\infty$,
where $\tau,\,g(\infty)$ are two constants independent of $z$ still to be determined.
@@
  • $y(z)=y(\infty)+\mathcal{O}(1/z)$, as $z\to\infty$,
where $y(\infty)$ is a constant still to be determined.
@@
$J_{C}(z)=\begin{cases}\begin{pmatrix}e^{\ln(N)\big{(}g_{-}(z)-g_{+}(z)\big{)}+y_{-}(z)-y_{+}(z)}&0\\ \\[2.0pt] i&e^{\ln(N)\big{(}g_{+}(z)-g_{-}(z)\big{)}+y_{+}(z)-y_{-}(z)}\end{pmatrix},&z\in\eta_{1}\\ \\[4.0pt] \begin{pmatrix}e^{\ln(N)\big{(}g_{-}(z)-g_{+}(z)\big{)}+y_{-}(z)-y_{+}(z)}&i\\ \\[2.0pt] 0&e^{\ln(N)\big{(}g_{+}(z)-g_{-}(z)\big{)}+y_{+}(z)-y_{-}(z)}\end{pmatrix},&z\in\eta_{2}\\ \\[4.0pt] \begin{pmatrix}e^{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{2}{\tau}}}\ln(N)-\Delta}&0\\ 0&e^{-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{2}{\tau}}}\ln(N)+\Delta}\end{pmatrix},&z\in\gamma\end{cases}$ (3.8)
@@ 3.1 The $g(z)$ function
$g^{\prime}(z)=-\frac{2c}{\tau R(z)},\quad R(z)=(z-A)^{\frac{1}{2}}(z+{\overline{A}})^{\frac{1}{2}}(z+A)^{\frac{1}{2}}(z-{\overline{A}})^{\frac{1}{2}}\,,$ (3.11)
for some $c,\tau\in{\mathbb{C}}$, where for each factor $(\cdot)^{1/2}$ we use the principal branch, so that $R(z)$ is analytic in $\mathbb{C}\setminus\{\eta_{1}\cup\eta_{2}$}, and is normalized such that
@@
$g_{+}(z)+g_{-}(z)=-\frac{4c}{\tau}\int_{\gamma}\frac{1}{R(s)}{\rm d}s+1,\quad z\in\eta_{2}$ (3.14)
$g_{+}(z)-g_{-}(z)=-\frac{4c}{\tau}\int_{\eta_{1}}\frac{1}{R_{+}(s)}{\rm d}s,\quad z\in\gamma$ (3.15)
Imposing
$c=\frac{1}{2}\left(\int_{\eta_{1}}\frac{1}{R_{+}(s)}{\rm d}s\right)^{-1}\,,\quad\tau=2c\int_{\gamma}\frac{1}{R(s)}{\rm d}s$ (3.16)
$g(z)$ in (3.12) solves RHP 3.3.
@@
  4. 4.
    $\int_{\eta_{1}}\frac{1}{R_{+}(z)}{\rm d}z\in{\mathbb{R}}_{-}$,
  5. 5.
    $\int_{\gamma}\frac{1}{R(z)}{\rm d}z\in i{\mathbb{R}}_{-}$.
@@
As a corollary of Proposition 3.6 we deduce that $c\in{\mathbb{R}}_{-}$ and $\tau\in i{\mathbb{R}}_{+}$, for $c,\tau$ defined in (3.16). So $\frac{2c}{\tau}\in i{\mathbb{R}}_{+}$.
@@ 3.2 The $y(z)$ function
$y(z)=\frac{R(z)}{2\pi i}\left(\int_{\eta_{1}}\frac{\log\left(2\pi h(s)\rho(s)\right){{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+2i\theta(s,x,t)}}}{R_{+}(s)(s-z)}{\rm d}s-\int_{\eta_{2}}\frac{\log\left(2\pi h(s)\rho(s)\right){{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2i\theta(s,x,t)}}}{R_{+}(s)(s-z)}{\rm d}s+\int_{\gamma}\frac{\Delta(x,t)}{R(s)(s-z)}{\rm d}s\right)\,,$ (3.18)
with $\Delta(x,t)$ defined as
$\Delta(x,t)=\left(\int_{\gamma}\frac{1}{R(s)}{\rm d}s\right)^{-1}\left(-\int_{\eta_{1}}\frac{\log\,(2\pi h(s)\rho(s))+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2i\theta(s,x,t)}}}{R_{+}(s)}{\rm d}s+\int_{\eta_{2}}\frac{\log(2\pi h(s)\rho(s))-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2i\theta(s,x,t)}}}{R_{+}(s)}{\rm d}s\right)\,.$ (3.19)
Furthermore, $\Delta(x,t)$ is independent of $t$, so that
$\Delta(x)=\left(\int_{\gamma}\frac{1}{R(s)}{\rm d}s\right)^{-1}\left(-\int_{\eta_{1}}\frac{\log\,(2\pi h(s)\rho(s))+2izx}{R_{+}(s)}{\rm d}s+\int_{\eta_{2}}\frac{\log\left(2\pi h(s)\rho(s)\right)-2ixz}{R_{+}(s)}{\rm d}s\right)\,.$ (3.20)
Lastly, $y(\infty)=\lim_{z\to\infty}y(z)$ is finite.
@@
The reader may verify using the Cauchy-Plemelj-Sokhotski formula that $y$ defined in (3.18) satisfies conditions (3.7).
@@
and (3.19) guarantees that $y(z)\to y_{\infty}$ as $z\to\infty$. Finally, we notice that $\Delta(x,t)$ is independent of time. Indeed, the coefficient of $t$ in $\Delta(x,t)$ is given by
$-4i\Bigg{(}\int_{\eta_{1}}\frac{s^{2}}{R_{+}(s)}{\rm d}s+\int_{\eta_{2}}\frac{s^{2}}{R_{+}(s)}{\rm d}s\Bigg{)}$ (3.22)
@@
$\Delta(x)=\left(\int_{\gamma}\frac{1}{R(s)}{\rm d}s\right)^{-1}\left(-\int_{\eta_{1}}\frac{\log\,(2\pi h(s)\rho(s))+2ixz}{R_{+}(s)}{\rm d}s+\int_{\eta_{2}}\frac{\log(2\pi h(s)\rho(s))-2ixz}{R_{+}(s)}{\rm d}s\right).$ (3.23)
For later use, we prove the following proposition.
Proposition 3.8.
Define $\Delta(x)$ as in (3.20), then $\Delta(x)\in i{\mathbb{R}}$.
Proof.
Applying Proposition 3.6, we know that $\int_{\gamma}\frac{1}{R(s)}{\rm d}s\in i{\mathbb{R}}$; so to conclude it remains to show that
$\left(-\int_{\eta_{1}}\frac{\log\,(2\pi h(s)\rho(s))+2ixz}{R_{+}(s)}{\rm d}s+\int_{\eta_{2}}\frac{\log\left(2\pi h(s)\rho(s)\right)-2ixz}{R_{+}(s)}{\rm d}s\right)\in{\mathbb{R}}\,.$ (3.24)
@@
$\begin{split}\int_{-a}^{a}\frac{\log\left(2\pi h({\widetilde{x}}-ib)\rho({\widetilde{x}}-ib)\right)-2ix({\widetilde{x}}-ib)}{R_{+}({\widetilde{x}}-ib)}{\rm d}{\widetilde{x}}&={\overline{\int_{-a}^{a}\frac{\log\left(2\pi h({\widetilde{x}}+ib)\rho({\widetilde{x}}+ib)\right)+2ix({\widetilde{x}}+ib)}{R_{+}({\widetilde{x}}+ib)}{\rm d}{\widetilde{x}}}}\\ &\stackrel{{\scriptstyle s={\widetilde{x}}+ib}}{{=}}{\overline{\int_{-{\overline{A}}}^{A}\frac{\log\left(2\pi h(s)\rho(s)\right)+2ixs}{R_{+}(s)}{\rm d}s}}\\ &=-{\overline{\int_{\eta_{1}}\frac{\log\left(2\pi h(s)\rho(s)\right)+2ixs}{R_{+}(s)}{\rm d}s}}\,,\end{split}$ (3.25)
from which (3.24) follows. ∎

### v2

@@ 2.5 Riemann-Hilbert problem for the soliton gas
where we recall that the contour $\eta_{1}$ is oriented from right to left, while $\eta_{2}$ is oriented from left to right, and in both cases the error term $\mathcal{O}\left(\frac{1}{N}\right)$ is uniform for all $z$ in the respective contour $\Gamma_{1}$ or $\Gamma_{2}$. Indeed, under the $N$-soliton condensate assumption 3, a complete asymptotic expansion for this sum can be calculated, because the eigenvalues $\{\lambda_{j}\}_{j=1}^{N}$ have been selected to arrive at a uniform mid-point approximation, which is known to converge with an error ${\mathcal{O}}(N^{-2})$. The next term in the expansion can be computed explicitly, we perform such computation in Appendix A.
@@
  • $\tilde{A}(z)=I+\frac{1}{2iz}\left(\begin{array}[]{cc}-\int_{x}^{\infty}|\psi_{SG}(s,t;N)|^{2}{\rm d}s&\psi_{SG}(x,t;N)\\ \overline{\psi_{SG}(x,t;N)}&\int_{x}^{\infty}|\psi_{SG}(s,t;N)|^{2}{\rm d}s\end{array}\right)+\mathcal{O}\left(\frac{1}{z^{2}}\right),\ \mbox{ as }z\to\infty$.
@@
This permits the direct application of an existence theory via the connection between Riemann-Hilbert problems and Fredholm singular integral operators (see Theorems 9.1, 9.2, and 9.3 in [43], or the proof of Theorem 5.3 in [11]). Regarding the asymptotic behaviour of the solution, using a standard dressing argument, see for example [4, 7], one can verify that the $z^{-1}$ coefficient in the large $z$ expansion for ${\widetilde{A}}(z)$ has the given form. Thus we state without further discussion the following result:
@@ 2.6 Statement of the results
Under the $N$-soliton condensate scattering data assumption, for all $(x,t)$ in a compact set ${\mathfrak{K}}$, there is $N_{0}$ so that for all $N>N_{0}$, $\psi_{SG}(x,t;N)$ satisfies the asymptotic description (7.7). Specifically, for $\psi_{SG}$, we have
$\begin{split}&\psi_{SG}(x,t;N)=\\ &-2i\frac{\Re(A)\Im(A)}{|A|}\operatorname{sd}\left(-2|A|x+\frac{2K(\cos(\theta))}{\pi}\ln(N)-\frac{2|A|}{\pi}\Re\left(\int_{\eta_{1}}\frac{\log(2\pi h(s)\rho(s))}{R_{+}(s)}{\rm d}s\right);\sin(\theta)\right)e^{it(A^{2}+{\overline{A}}^{2})+i\phi_{0}}\\ &+\mathcal{O}\left(\frac{1}{\log{N}}\right),\end{split}$ (2.35)
where the error term $\mathcal{O}\left(\frac{1}{\log{N}}\right)$ is uniform for all $(x,t)$ in the compact set ${\mathfrak{K}}$. Here $\operatorname{sd}$ is the Jacobi elliptic function $\operatorname{sd}$ [13, Chap. 22], and $R(z)$ and $\phi_{0}$ are given by
$R(z)=(z-A)^{\frac{1}{2}}(z+A)^{\frac{1}{2}}(z-{\overline{A}})^{\frac{1}{2}}(z+{\overline{A}})^{\frac{1}{2}}\,,$ (2.36)
$\phi_{0}=\ln(N)\frac{K(\cos(\theta))}{K(\sin(\theta))}+\Re\left(\int_{\eta_{1}}\frac{s\ln(2\pi h(s)\rho(s))}{R_{+}(s)}{\rm d}s\right)\,,\quad\theta=\arg(A)$ (2.37)
where the function $K(k)$ is the complete elliptic integral of the fist kind [13, Chap. 19], the branch-cuts are the canonical ones, and the integration is performed on the straight line connecting the endpoints.
@@
We notice that, in view of the Galilean invariance of the NLS equation, for $v\in{\mathbb{R}}$, $\psi(x-vt,t)e^{ivx-i\frac{v^{2}t}{2}}$ solves the NLS equations (1.1) whenever $\psi(x,t)$ does. Furthermore, if the scattering maps gives $\psi(x,0)\to\{r(z),\{\lambda_{j},c_{j}\}_{j=1}^{N}\}$ then $\psi(x,0)e^{ivx}\to\{r\left(z+\frac{v}{2}\right),\{\lambda_{j}-\frac{v}{2},c_{j}\}_{j=1}^{N}\}$; therefore, in the setting where the two line segments $\eta_{1}$ and $\eta_{2}$ are shifted so that they are not symmetric with respect to the origin, one can always reduce to the symmetric situation up to a simple scaling. Finally, it is worth noticing that the solution depends just on the product of $h(z),\rho(z)$ and not on each function separately. Fig. 3 shows plots of the leading order asymptotic of the solution of the NLS equation for different values of $N$.
@@
Let $A_{N}(z),{\widetilde{A}}(z)$ be the solutions of RHP 2.3 and RHP 2.4 respectively. Under the $N$-soliton condensate scattering data assumption, there is $N_{0}$ so that for all $N>N_{0}$, we have
$A_{N}(z)=\left(I+\mathcal{O}\left(\frac{1}{N\left(1+|z|\right)}\right)\right){\widetilde{A}}(z)\,.$ (2.38)
where the error term, $\mathcal{O}\left(\frac{1}{N\left(1+|z|\right)}\right)$ is uniform in the entire complex plane, and $(x,t)$ in a compact set.
@@
In the next four sections, we develop the asymptotic analysis of the solution of the Riemann-Hilbert problem 2.4, to establish, in Section 7, the asymptotic behavior of the soliton gas condensate solution $\psi_{SG}(x,t;N)$. In Section 8 we consider a more complex solution which contains one additional soliton, and using our analysis we provide a rigorous validation of the kinetic theory of solitons in this setting. We prove Proposition 2.7, Corollary 2.8 and other technical results in Appendix A.
@@ 3 Dressing transformations
$C(z)=e^{(\ln(N)g_{\infty}+y_{\infty})\sigma_{3}}B(z)e^{-(\ln(N)g(z)+y(z))\sigma_{3}}\,,$ (3.3)
where $g_{\infty},y_{\infty}$ are the values of $g(z)$ and $y(z)$ at $\infty$, which we assume are finite (this assumption will be verified later, since these two quantities will be explicitly constructed). The function $C(z)$ satisfies the following RHP.
@@
    $g_{+}(z)-g_{-}(z)=c_{1},\quad z\in\gamma$ (3.6)
  • $g(z)=g_{\infty}+\mathcal{O}(1/z)$, as $z\to\infty$,
where $c_{1},\,g_{\infty}$ are two constants independent of $z$ still to be determined.
@@
  • $y(z)=y_{\infty}+\mathcal{O}(1/z)$, as $z\to\infty$,
where $y_{\infty}$ is a constant still to be determined.
@@
$J_{C}(z)=\begin{cases}\begin{pmatrix}e^{\ln(N)\big{(}g_{-}(z)-g_{+}(z)\big{)}+y_{-}(z)-y_{+}(z)}&0\\ \\[2.0pt] i&e^{\ln(N)\big{(}g_{+}(z)-g_{-}(z)\big{)}+y_{+}(z)-y_{-}(z)}\end{pmatrix},&z\in\eta_{1}\\ \\[4.0pt] \begin{pmatrix}e^{\ln(N)\big{(}g_{-}(z)-g_{+}(z)\big{)}+y_{-}(z)-y_{+}(z)}&i\\ \\[2.0pt] 0&e^{\ln(N)\big{(}g_{+}(z)-g_{-}(z)\big{)}+y_{+}(z)-y_{-}(z)}\end{pmatrix},&z\in\eta_{2}\\ \\[4.0pt] \begin{pmatrix}e^{-c_{1}\ln(N)-\Delta}&0\\ 0&e^{c_{1}\ln(N)+\Delta}\end{pmatrix},&z\in\gamma\end{cases}$ (3.8)
@@ 3.1 The $g(z)$ function
$g^{\prime}(z)=\frac{c_{0}}{R(z)},\quad R(z)=(z-A)^{\frac{1}{2}}(z+{\overline{A}})^{\frac{1}{2}}(z+A)^{\frac{1}{2}}(z-{\overline{A}})^{\frac{1}{2}}\,,$ (3.11)
for some $c_{0}\in{\mathbb{C}}$, where for each factor $(\cdot)^{1/2}$ we use the principal branch, so that $R(z)$ is analytic in $\mathbb{C}\setminus\{\eta_{1}\cup\eta_{2}$}, and is normalized such that
@@
$g_{+}(z)+g_{-}(z)=2c_{0}\int_{\gamma}\frac{1}{R(s)}{\rm d}s+1,\quad z\in\eta_{2}$ (3.14)
$g_{+}(z)-g_{-}(z)=2c_{0}\int_{\eta_{1}}\frac{1}{R_{+}(s)}{\rm d}s,\quad z\in\gamma$ (3.15)
Imposing the conditions
$c_{0}=-\left(\int_{\gamma}\frac{1}{R(s)}{\rm d}s\right)^{-1}\,\mbox{and}\quad c_{1}=2c_{0}\int_{\eta_{1}}\frac{1}{R(s)}{\rm d}s\,,$ (3.16)
one may verify that $g(z)$ in (3.12) solves RHP 3.3.
@@
Furthermore,
$\int_{\eta_{1}}\frac{1}{R_{+}(z)}{\rm d}z=-\frac{K(\cos(\theta))}{|A|}\,,\quad\int_{-{\overline{A}}}^{-A}\frac{1}{R(s)}{\rm d}s=-i\frac{K(\sin(\theta))}{|A|}\,,\,,$ (3.18)
where $\theta=\arg(A)$, and $K(k)$ is the complete elliptic integral of the first kind [13, Chap. 19].
@@
For later usage, we define
$c=\frac{1}{2\int_{\eta_{1}}\frac{1}{R_{+}(z)}{\rm d}z}=-\frac{|A|}{2K(\cos(\theta))}\,,\quad\tau=\frac{\int_{-{\overline{A}}}^{-A}\frac{1}{R(z)}{\rm d}z}{\int_{\eta_{1}}\frac{1}{R_{+}(z)}{\rm d}z}=2i\frac{K(\cos(\theta))}{K(\sin(\theta))}\,,$ (3.19)
and we notice that $c\in{\mathbb{R}}_{-}$, $\tau\in i{\mathbb{R}}_{+}$, and $\frac{2c}{\tau}\in i{\mathbb{R}}_{+}$. In this notation
$c_{0}=-\frac{2c}{\tau}\,,\qquad c_{1}=-\frac{2}{\tau}\,,$ (3.20)
therefore, one can rewrite $g(z)$ as
$g(z)=-\frac{2c}{\tau}\int_{A}^{z}\frac{{\rm d}s}{R(s)}+\frac{1}{2}$ (3.21)
Furthermore, as a corollary of Proposition 3.6 the following holds:
$\frac{2c}{\tau}\int_{-{\overline{A}}}^{-A}\frac{1}{R(s)}{\rm d}s=1,\quad\frac{2c}{\tau}\int_{A}^{-{\overline{A}}}\frac{1}{R_{+}(s)}{\rm d}s=\frac{1}{\tau},\quad-\frac{2c}{\tau}\int_{A}^{\infty}\frac{1}{R(s)}{\rm d}s=\frac{1}{2\tau}-\frac{1}{2}\,.$ (3.22)
@@ 3.2 The $y(z)$ function
$y(z)=\frac{R(z)}{2\pi i}\left(\int_{\eta_{1}}\frac{\log\left(2\pi h(s)\rho(s)\right){{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+2i\theta(s,x,t)}}}{R_{+}(s)(s-z)}{\rm d}s-\int_{\eta_{2}}\frac{\log\left(2\pi h(s)\rho(s)\right){{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2i\theta(s,x,t)}}}{R_{+}(s)(s-z)}{\rm d}s+\int_{\gamma}\frac{\Delta(x)}{R(s)(s-z)}{\rm d}s\right)\,,$ (3.23)
with $\Delta(x)$
$\Delta(x)=-\frac{2c}{\tau}\left(2\pi x+2\Re\left(\int_{\eta_{1}}\frac{\log\,(2\pi h(s)\rho(s))}{R_{+}(s)}{\rm d}s\right)\right)\,.$ (3.24)
Lastly, $y_{\infty}=\lim_{z\to\infty}y(z)$ is finite.
Remark 3.2.
We notice that $\Delta(x)\in i{\mathbb{R}}$.
@@
By applying the Cauchy-Plemelj-Sokhotski formula we obtain that the function $y(z)$ has the following form
$y(z)=\frac{R(z)}{2\pi i}\left(\int_{\eta_{1}}\frac{\log\left(2\pi h(s)\rho(s)\right){{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+2i\theta(s,x,t)}}}{R_{+}(s)(s-z)}{\rm d}s-\int_{\eta_{2}}\frac{\log\left(2\pi h(s)\rho(s)\right){{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2i\theta(s,x,t)}}}{R_{+}(s)(s-z)}{\rm d}s+\int_{\gamma}\frac{{\widehat{\Delta}}(x,t)}{R(s)(s-z)}{\rm d}s\right)\,,$ (3.25)
with ${\widehat{\Delta}}(x,t)$ defined as
${\widehat{\Delta}}(x,t)=\frac{2c}{\tau}\left(-\int_{\eta_{1}}\frac{\log\,(2\pi h(s)\rho(s))+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2i\theta(s,x,t)}}}{R_{+}(s)}{\rm d}s+\int_{\eta_{2}}\frac{\log(2\pi h(s)\rho(s))-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2i\theta(s,x,t)}}}{R_{+}(s)}{\rm d}s\right)\,.$ (3.26)
@@
and (3.26) guarantees that $y_{\infty}:=\lim_{z\to\infty}y(z)$ exists. To conclude, we must show that ${\widehat{\Delta}}(x,t)=\Delta(x)$ in (3.24). First, we notice that ${\widehat{\Delta}}(x,t)$ is independent of time. Indeed, the coefficient of $t$ in ${\widehat{\Delta}}(x,t)$ is given by
$-2i\Bigg{(}\int_{\eta_{1}}\frac{s^{2}}{R_{+}(s)}{\rm d}s+\int_{\eta_{2}}\frac{s^{2}}{R_{+}(s)}{\rm d}s\Bigg{)}$ (3.28)
@@
${\widehat{\Delta}}(x,t)=\left(\int_{\gamma}\frac{1}{R(s)}{\rm d}s\right)^{-1}\left(-\int_{\eta_{1}}\frac{\log\,(2\pi h(s)\rho(s))+2ixs}{R_{+}(s)}{\rm d}s+\int_{\eta_{2}}\frac{\log(2\pi h(s)\rho(s))-2ixs}{R_{+}(s)}{\rm d}s\right).$ (3.29)
Moreover, by residue calculation, one can show that
$-2ix\left(\int_{\eta_{1}}\frac{s}{R(s)}{\rm d}s+\int_{\eta_{1}}\frac{s}{R(s)}{\rm d}s\right)=-2\pi x\,.$ (3.30)
Finally, applying Proposition 3.6, we know that $\int_{\gamma}\frac{1}{R(s)}{\rm d}s\in i{\mathbb{R}}$; so to conclude it remains to show that
$\left(-\int_{\eta_{1}}\frac{\log\,(2\pi h(s)\rho(s))}{R_{+}(s)}{\rm d}s+\int_{\eta_{2}}\frac{\log\left(2\pi h(s)\rho(s)\right)}{R_{+}(s)}{\rm d}s\right)=-2\Re\left(\int_{\eta_{1}}\frac{\log\,(2\pi h(s)\rho(s))}{R_{+}(s)}{\rm d}s\right)\,.$ (3.31)
@@
$\begin{split}\int_{-a}^{a}\frac{\log\left(2\pi h({\widetilde{x}}-ib)\rho({\widetilde{x}}-ib)\right)}{R_{+}({\widetilde{x}}-ib)}{\rm d}{\widetilde{x}}&={\overline{\int_{-a}^{a}\frac{\log\left(2\pi h({\widetilde{x}}+ib)\rho({\widetilde{x}}+ib)\right)}{R_{+}({\widetilde{x}}+ib)}{\rm d}{\widetilde{x}}}}\\ &\stackrel{{\scriptstyle s={\widetilde{x}}+ib}}{{=}}{\overline{\int_{-{\overline{A}}}^{A}\frac{\log\left(2\pi h(s)\rho(s)\right)}{R_{+}(s)}{\rm d}s}}\\ &=-{\overline{\int_{\eta_{1}}\frac{\log\left(2\pi h(s)\rho(s)\right)}{R_{+}(s)}{\rm d}s}}\,,\end{split}$ (3.32)
from which it follows that ${\widehat{\Delta}}(x,t)\equiv\Delta(x)$, which completes the proof. ∎